Water leaves a hose at a rate of 1.5 kg/s with a speed of 20 m/s and is aimed at the side of a car which stops it. Assuming no splashing back, what is the force exerted by the water?
1
Expert's answer
2012-05-09T07:57:37-0400
The rate of change of momentum is equal to the net force:

F = dp/dt.

This can be shown using Newton’s second law. So,

F = ΔP/Δt = (P_final - P_initial)/Δt = (0 - 30kg·m/s) / 1.0s = -30N.